c) Watamu Ltd expects to pay a dividend of sh.7 per each ordinary share at the end of the year. Dividends are expected to grow at a rate of 6% per annum. The company’s policy is to retain 30% of its earnings and the return expected by the shareholders is 10%. Required (i) The intrinsic value of Watamu Ltd ordinary share (5 marks) (ii) The earning per share of the company (2 marks)
Answer #1

(i) calculation of intrinsic value of a share of Watamu Ltd :

formula :

where the dividend per share and constant growth rate are given

Intrinsic value = dividend per share / (cost of capital - growth rate)

= 7 / ( 10 % - 6 %)

= 7 / 4%

= 7 / 0.04

= Sh. 175

The intrinsic value per share is Sh. 175

(ii) The earnings per share of Watamu ltd is :

formula:

dividend payout = dividend per share / earnings per share

so from above formula

earnings per share = dividend per share / dividend pay out

= 7 / 70 %

= 7 / 0.70

= Sh.10

the earnings per share is Sh. 10

* for the calculation of dividend payout % refer working note :

working note:-

dividend pay out = 100% - retained earnings %

= 100 % - 30 %

= 70 %